OpenStudy (anonymous):

Find the exact value of each trig function WITHOUT a calculator. sin(x-y) if sin x=7/25, cos x=24/25, sin y=3/5 and cos y=4/5

OpenStudy (anonymous):

\[ \text{Note: } \sin(a - b) \equiv \sin(a)\cos(b) - \sin(b)\cos(a) \\ \text{Therefore: } \sin(x - y) \equiv \sin(x)\cos(y) - \sin(y)\cos(x) \\ \text{Which becomes: } \frac{7}{25} * \frac{4}{5} - \frac{3}{5} * \frac{24}{25} \equiv -\frac{44}{125}\]